OpenStudy (bossimbacon):

which equation shows the substitution method being used to solve the system of linear equations? x - y= 7 x = y + 3 @pooja195

pooja195 (pooja195):

Where are the equations ?

OpenStudy (bossimbacon):

A) x - ( y + 3) = 7

OpenStudy (bossimbacon):

B) ( y + 3) - y= 7

OpenStudy (bossimbacon):

C) x =(7 - x) + 3

OpenStudy (bossimbacon):

D) x - y = y + 3

pooja195 (pooja195):

What do you think?

OpenStudy (bossimbacon):

i dont really know

pooja195 (pooja195):

Take equation 2 and plug it into equation 1 in place of x since it gives you what x ='s

OpenStudy (bossimbacon):

i have a question about that, how do i know which equation to start with to plug into the other one, is it always the second one or??

OpenStudy (bossimbacon):

so it would b (B) then

pooja195 (pooja195):

Correct

pooja195 (pooja195):

And you would know what equation to start with if you see that a vairble is alone (if it is not alone isolate it)

OpenStudy (bossimbacon):

thank you so much!

pooja195 (pooja195):

yw ^_^